OpenStudy (anonymous):

For which real values of p does the series ∞∑(n=1) n=π*(ln p)^n converge? State reasons for your answer

OpenStudy (anonymous):

This? \[\sum_{j=1}^{\infty} \pi \ln^j(p)?\]

OpenStudy (jamesj):

Or this? \[\sum_{n=1}^{\infty} \pi (\ln p)^n\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

The second one.

OpenStudy (jamesj):

Wel, this is just a geometric series in disguise: a, ar, ar^2, ar^3, ... When does the sum of a geometric series converge and what are a and r here?

OpenStudy (anonymous):

Ah yep. Converges when r < 1, a = π, r = (lnp)^n ?

OpenStudy (zarkon):

you are a little off there

OpenStudy (anonymous):

r=ln(p) and what we both wrote for the series is the exact same. Just different notation.

OpenStudy (anonymous):

Since ln(1/e)=-1 and the ln(e)=1 then it has to be: 1/e<p<e

OpenStudy (jamesj):

Yes The infinite sum of a geometric series (ar^j) converges if and only if |r| < 1

OpenStudy (anonymous):

Okay awesome. Thanks for all the help everyone.